# TabNSA: Native Sparse Attention for Efficient Tabular Data Learning thanks: Code on GitHub: https://github.com/aseslamian/TabNSA

### v1

@@ Abstract
Tabular data poses unique challenges for deep learning due to its heterogeneous features and lack of inherent spatial structure. This paper introduces TabNSA, a novel deep learning architecture leveraging Native Sparse Attention (NSA) specifically for efficient tabular data processing. TabNSA incorporates a dynamic hierarchical sparse strategy, combining coarse-grained feature compression with fine-grained feature selection to preserve both global context awareness and local precision. By dynamically focusing on relevant subsets of features, TabNSA effectively captures intricate feature interactions. Extensive experiments demonstrate that TabNSA consistently outperforms existing methods, including both deep learning architectures and ensemble decision trees, achieving state-of-the-art performance across various benchmark datasets.
Index Terms:
Tabular data, Native Sparse Attention
I Introduction
Tabular data, ubiquitous across diverse domains, often presents challenges for traditional deep learning models due to its heterogeneous nature and lack of inherent spatial structure. While deep learning has revolutionized fields like computer vision and natural language processing, its application to tabular data has been less successful, with ensemble decision trees continuing to dominate many real-world applications [1, 2]. This discrepancy stems from the fact that canonical architectures designed for encoding raw data into meaningful representations have not yet been effectively adapted for tabular data [3]. The relatively inferior performance of deep learning models on tabular data can be attributed to factors such as limited labeled training data and the models’ struggle to capture complex relationships [4]. Researchers are actively exploring innovative techniques to address the limitations of decision trees and expand the applicability of deep learning in this domain [5].
This paper introduces TabNSA, inspired by Natively Sparse Attention (NSA) [6], which adapts the sparse attention mechanism for tabular data. TabNSA employs dynamic sparse attention to process heterogeneous features efficiently.
Instead of treating all features uniformly, it learns to focus on the most relevant subset for each instance, dynamically adjusting feature importance. This hierarchical sparse attention allows TabNSA to capture both global context and fine-grained feature dependencies, performing coarse selection of the most informative features and then modeling their interactions in detail.
The key contributions of this work are as follows:
  • TabNSA Architecture: We propose TabNSA, a novel deep learning model that integrates Native Sparse Attention into tabular data learning. By dynamically selecting and attending to relevant subsets of features, TabNSA effectively addresses the challenge of feature heterogeneity and captures complex feature dependencies. This approach provides a strong foundation for multimodal applications.
  • State-of-the-Art Performance: Our comprehensive experiments demonstrate that TabNSA consistently outperforms existing methods, including both deep learning architectures and ensemble decision trees. TabNSA achieves state-of-the-art accuracy on most benchmark datasets, demonstrating its strong generalization capabilities across diverse tabular tasks.
II Related Work
@@
Existing tabular learning methods aim to address various challenges associated with diverse datasets. These include situations where the feature sets differ between training and test data, limited or missing labels, and the potential for new features to emerge over time [7]. Classical techniques include parametric methods like Logistic Regression and non-parametric approaches such as K-Nearest Neighbors, Gradient Boosting, and Decision Trees [8]. Various models have been developed, including logistic regression, XGBoost [9, 10], and Multi-Layer Perceptrons.
@@
III Background
To address the computational inefficiency of standard attention mechanisms in long-context modeling for tabular data processing, we adapt Native Sparse Attention—a natively trainable sparse attention architecture [18, 19]. This adaptation tackles the quadratic complexity of standard attention mechanisms by leveraging a hierarchical sparse strategy that combines token compression, blockwise token selection, and local sliding windows. Our goal is to enhance computational efficiency while maintaining both global and local context awareness in tabular data applications.
In the model, each sample’s feature vector is first embedded and then projected into queries, keys, and values:
@@ III Background
    After the sparse attention module combines the outputs from the different branches, the resulting representation is processed by further linear layers (and activation functions like GELU) to produce the final regression output. The final output is as follows:
    $o_{t}^{*}=\sum_{c\in\mathcal{C}}g_{t}^{c}\cdot\text{Attn}(q_{t},\tilde{K}_{t}^{c},\tilde{V}_{t}^{c})$ (10)
    $\hat{y}=f(o_{t}^{*}).$ (11)
@@
    The gate score, $g^{c}_{t}\in[0,1]$, is obtained from the input features using sigmoid activation and an MLP and $f$ is final linear layer.
For tabular data learning, NSA can be adapted by treating each feature as a token and learning the relationships between these features through its sparse attention mechanisms. This approach is particularly beneficial for high-dimensional tabular data, where the computationally intensive standard attention mechanisms would be prohibitive.
IV Experiments and Results
IV-A Dataset
@@
In this work, we evaluated the performance of the TabNSA model and several baseline models across a diverse set of datasets. The assessment included 8 widely used binary classification datasets, 2 multi-class datasets, and 3 regression datasets sourced from the UCI repository, OpenML, and Kaggle. These datasets span a wide range of domains, including finance, business, chemistry, geography, ecology, image recognition, and sports, ensuring a comprehensive evaluation of the models. This diverse benchmark dataset collection enables us to systematically assess the performance across various domains and data structures. A detailed description of each dataset is provided in Table I.
@@
TABLE I: Dataset details including abbreviation, number of classes, number of data points, and number of features.
@@
Sarcos-Robotics SA - 48,933 28
Topo-2-1 TO - 8,885 266
CPU-Predict CP - 8,192 13
IV-B Preprocessing

### v2

@@ Abstract
Tabular data poses unique challenges for deep learning due to its heterogeneous feature types, lack of spatial structure, and often limited sample sizes. We propose TabNSA, a novel deep learning framework that integrates Native Sparse Attention (NSA) with a TabMixer backbone to efficiently model tabular data. TabNSA tackles computational and representational challenges by dynamically focusing on relevant feature subsets per instance. The NSA module employs a hierarchical sparse attention mechanism, including token compression, selective preservation, and localized sliding windows, to significantly reduce the quadratic complexity of standard attention operations while addressing feature heterogeneity. Complementing this, the TabMixer backbone captures complex, non-linear dependencies through parallel multilayer perceptron (MLP) branches with independent parameters. These modules are synergistically combined via element-wise summation and mean pooling, enabling TabNSA to model both global context and fine-grained interactions. Extensive experiments across supervised and transfer learning settings show that TabNSA consistently outperforms state-of-the-art deep learning models. Furthermore, by augmenting TabNSA with a fine-tuned large language model (LLM), we enable it to effectively address Few-Shot Learning challenges through language-guided generalization on diverse tabular benchmarks.
1 Introduction
Tabular data, prevalent in domains like finance, healthcare, and e-commerce, poses unique challenges due to its heterogeneity, lack of spatial or temporal structure, non-uniform feature interactions, and often limited sample sizes [1]. Traditional models like Gradient Boosted Decision Trees (GBDT) [2, 3] excel in these settings, but deep learning models often struggle to match their performance.
In response to these challenges, researchers have proposed a variety of deep learning models tailored to tabular data. Some approaches adapt multilayer perceptrons (MLPs) with enhanced regularization or embedding mechanisms, while others introduce architectural innovations such as attentive feature selection, gating mechanisms, or sparsity-inducing modules. Notable examples include TabNet [4], which utilizes attentive feature masks to model sparse feature dependencies, and DSelect-k [5], which provides a differentiable relaxation of feature subset selection. More recently, Transformer-inspired models have been explored for capturing pairwise and higher-order feature interactions [6]. While these methods have advanced the field, many remain computationally expensive, lack dynamic sparsity, or fail to generalize across tasks with varying data regimes, such as low-data or feature-incremental settings.
To bridge this gap, recent research has proposed deep learning architectures specifically tailored for tabular data. One such model, TabMixer [7], extends the MLP-Mixer architecture by employing channel-wise and token-wise mixing to capture global feature and instance dependencies. It achieves strong performance on supervised, transfer and incremental learning tasks while maintaining low computational complexity and robustness to missing values. However, a key limitation of TabMixer is that it processes all features uniformly, without distinguishing instance-specific feature importance, which potentially leads to redundancy and reduced interpretability.
In parallel, Native Sparse Attention (NSA) [8] has demonstrated the effectiveness of sparse attention mechanisms in reducing computational overhead while preserving performance in sequence modeling. Inspired by this, we propose TabNSA, a new deep learning framework that integrates NSA with the TabMixer backbone to address both computational and modeling challenges in tabular learning.
TabNSA introduces a hierarchical sparse attention mechanism that dynamically selects a relevant subset of features for each instance, effectively reducing redundancy and highlighting instance-specific dependencies. This coarse-to-fine selection is followed by TabMixer’s MLP-based token and channel mixing, which efficiently models the complex interactions of the selected features. NSA enhances the model adaptability and scalability, while TabMixer provides expressive capacity to capture non-linear dependencies. Their integration results in a lightweight yet powerful architecture capable of handling high-dimensional, heterogeneous tabular data.
Adapting NSA to tabular data involves treating features as tokens, shifting the attention paradigm from sequences (e.g., time steps or words) to feature dimensions. This design choice is motivated by the observation that in tabular datasets, not all features are equally informative for every instance, and their interactions are often sparse and heterogeneous. Treating features as tokens allows the model to dynamically allocate attention to a subset of relevant dimensions, enabling instance-specific modeling of feature relevance. The hierarchical sparse attention of the NSA, comprising token compression, selection, and localized attention, naturally aligns with these properties. In particular, the token compression mechanism mitigates redundancy among correlated features, while selective preservation allows the model to focus computational resources on informative features, addressing both heterogeneity and sample efficiency. Compared to standard dense attention, which treats all features uniformly and scales quadratically with feature count, NSA provides a principled and scalable alternative that is better suited for high-dimensional, noisy tabular inputs. In our ablation studies (see Experiment section), we empirically confirm that NSA outperforms full-attention variants in both accuracy and efficiency across diverse tabular benchmarks.
In summary, this paper makes the following contributions:
  • TabNSA Architecture: We propose TabNSA, a novel deep learning model integrating NSA with a TabMixer-based representation to dynamically select and process instance-specific feature subsets. This fusion effectively captures both global and local feature interactions, enhancing efficiency and expressiveness for tabular data.
  • Efficient Hierarchical Sparse Attention: TabNSA employs effective hierarchical sparse attention mechanisms, significantly reducing attention complexity without sacrificing performance. This enables the model to scale efficiently to large tabular datasets.
  • Efficiency and Generalization: By leveraging sparse attention and hierarchical feature processing, TabNSA efficiently captures long-range dependencies and feature-wise interactions. This contributes to strong generalization across various tasks and facilitates insightful feature importance analysis, enhancing interpretability.
  • Enhanced Few-Shot and Transfer Learning Capabilities: Through integration with large language models such as Gemma, TabNSA demonstrates superior performance in few-shot and low-resource learning scenarios, significantly outperforming existing baselines.
  • State-of-the-Art Performance: Extensive empirical evaluation across supervised, transfer, and few-shot learning tasks demonstrates that TabNSA consistently achieves state-of-the-art performance, surpassing leading deep learning architectures, particularly in low-resource and transfer settings.
To situate TabNSA within the broader literature, we next review related work in tabular learning and sparse attention mechanisms.
2 Related Work
@@
Existing tabular learning methods aim to address various challenges associated with diverse datasets. These include situations where the feature sets differ between training and test data, limited or missing labels, and the potential for new features to emerge over time [9]. We will briefly review related work in the following.
Classical techniques include parametric methods like Logistic Regression and non-parametric approaches such as K-Nearest Neighbors, Gradient Boosting, and Decision Trees [10]. Various models have been developed, including logistic regression, XGBoost [11, 12], and Multi-Layer Perceptrons.
@@
Sparse Attention Mechanisms. Sparse attention has emerged as a powerful paradigm for improving the efficiency and interpretability of deep models, particularly in high-dimensional settings. In NLP, methods such as Longformer, BigBird, and Routing Transformer [19, 20, 21] introduce sparsity into the attention matrix to handle long sequences more efficiently without sacrificing contextual understanding. In computer vision, sparse attention has been used to focus on salient spatial regions while reducing computation, as seen in models like SparseViT and TokenLearner [22]. The recent Natively Sparse Attention (NSA) framework [8] extends this idea by learning instance-specific sparse attention patterns through token compression, selection, and local context modeling, achieving competitive performance with reduced cost in language and vision tasks. Unlike fixed sparse patterns (e.g., in BigBird), NSA learns data-driven sparsity patterns per instance, which is critical in tabular data where relevant features vary
We draw inspiration from these advances and adapt NSA to tabular data, where high dimensionality, feature redundancy, and heterogeneous feature relevance across instances present similar challenges. By treating features as tokens, sparse attention enables TabNSA to dynamically select a subset of informative features per instance—mirroring the instance-adaptive attention behavior that proved successful in sequence and image domains. This adaptation helps improve computational scalability and enhances interpretability for tabular learning tasks.
3 Background
To address the computational inefficiency of standard attention mechanisms in long-context modeling for tabular data processing, we adapt Native Sparse Attention, a natively trainable sparse attention architecture [23, 24]. Additionally, we incorporate TabMixer [7], which captures complex patterns and global dependencies inherent in tabular data, where feature relationships are often non-linear and interdependent.
3.1 Native Sparse Attention
For tabular data learning, NSA can be adapted by treating each feature as a token and learning the relationships between these features through its sparse attention mechanisms. In the architecture, each sample’s feature vector is first embedded and then projected into queries, keys, and values:
@@ 3.1 Native Sparse Attention
    After the sparse attention module combines the outputs from the different branches, the resulting representation is processed by further linear layers (and activation functions like GELU) to produce the final output as follows:
    $Y=\sum_{c\in\mathcal{C}}g_{t}^{c}\cdot\text{Attn}(q_{t},\tilde{K}_{t}^{c},\tilde{V}_{t}^{c})$ (10)
@@
    The gate score, $g^{c}_{t}\in[0,1]$, is obtained from the input features using sigmoid activation and an MLP.
3.2 TabMixer
The TabMixer block [7] is an enhanced version of the standard MLP-Mixer block, designed to process information through three main branches: a residual connection, a central pathway for channel-wise mixing, and a parallel path for token-wise mixing.
Unlike the standard MLP-Mixer which uses tied parameters in its MLPs, the TabMixer block utilizes parallel MLPs with independent learning parameters. This parallel structure, combined with an additional SiLU non-linear activation function, allows the model to more effectively capture complex, non-linear, and global dependencies within tabular data by exploring interactions across both features and examples simultaneously.
The block’s operation can be described mathematically. For an input matrix $X\in\mathbb{R}^{B\times N}$ the output Z is computed as follows:
$Z=\text{SiLU}\left[\text{GeLU}\left(\text{MLP}_{1}(X^{T})\right)^{T}\odot\text{MLP}_{2}(X)\right]+X$ (11)
$\text{MLP}_{1,2}(X)=W_{1,2}\text{LayerNorm}(X)+b_{1,2}$
The final output produces attention-aware features through additive adjustments.
3.3 Gemma
Gemma [25] is a lightweight, open-source family of large language models (LLMs) developed by Google DeepMind, leveraging the technology behind Gemini models [26]. Built upon a decoder-only Transformer architecture, Gemma models are designed for computational efficiency while maintaining strong performance across various natural language tasks. The Gemma-1.1-2B-IT variant, specifically employed in our study, features 2 billion parameters and has undergone instruction tuning. Instruction tuning enhances the model’s capability to interpret natural language prompts, making it particularly effective for zero-shot or few-shot learning scenarios[27] [28]. This capability is instrumental in our approach, where structured tabular data is transformed into textual prompts, harnessing Gemma’s advanced text representation and reasoning abilities for encoding tasks. Gemma’s recent enhancements include extended context lengths, multilingual support, and a novel local-global attention mechanism for memory efficiency.
We selected the Gemma family of LLMs due to its strong performance on open-ended reasoning tasks and its flexibility in adapting to various downstream applications with minimal fine-tuning. Unlike encoder-only models such as BERT or RoBERTa, which are optimized for classification and masked language modeling, Gemma follows a decoder-style architecture more aligned with instruction tuning and generative few-shot prompting. This structure allows Gemma to better leverage natural language context and examples when guiding few-shot learning. In our preliminary tests, Gemma outperformed BERT and RoBERTa variants on few-shot tabular classification when using text-formatted input prompts, especially in extremely low-data regimes.
In our methodology, we freeze most encoder parameters of the pretrained Gemma to preserve generalized linguistic features and selectively fine-tune only the last Transformer layers. The resultant embeddings undergo a dual-pooling and linear-adaptation process before integration into our TabNSA classification module. This combined approach leverages Gemma’s pretrained linguistic power to ensure robust generalization and efficiency in few-shot learning contexts. In the next section, we describe how NSA’s sparse mechanisms address feature redundancy and TabMixer’s nonlinear mixing captures complex interactions, forming TabNSA.
4 Method
4.1 Adapting NSA to Tablular Data
First, we have adapted vanilla NSA to tabular data. The adaptation is detailed in Table 1, which illustrates the conceptual mapping between the original use of NSA in sequence modeling and its adaptation for feature-wise modeling in tabular data.
Aspect NSA in Sequence Modeling NSA in Tabular Data (TabNSA)
Token definition Each token represents a time step or word in a sequence. Each token corresponds to a feature (i.e., a column in the table).
Input structure Ordered sequence of elements (e.g., text or time series). Unordered, heterogeneous set of features.
Attention focus Focuses on relevant past or future tokens. Focuses on the most informative features per instance.
Token compression Reduces temporal redundancy (e.g., repeated words or uninformative frames). Reduces feature redundancy (e.g., correlated or irrelevant features).
Local sliding window Preserves local temporal patterns. Captures local dependencies among related feature groups.
Sparse selection effect Enables efficient modeling of long sequences. Enables efficient learning on high-dimensional, sparse tabular inputs.
Interpretability Highlights influential time points or tokens. Highlights per-instance key features, aiding interpretation.
Table 1: Conceptual adaptation of NSA from sequence modeling to tabular data modeling.
4.2 TabNSA Model Description
Next, to effectively capture both long-range dependencies and feature-wise interactions in tabular data, we introduce TabNSA, a hybrid neural architecture that integrates Native Sparse Attention (NSA) with TabMixer. The model is designed to efficiently process tabular inputs by leveraging sparse self-attention for context-aware representation and TabMixer for non-linear, cross-feature interaction modeling.
[…]
Figure 2: TabMixer (Right) and NSA (Left) Modules
Given an input tensor $X\in\mathbb{R}^{B\times N}$, where $B$ is the batch size and $N$ is the number of features, the model performs the following steps:
  1. 1.
    Feature Embedding: Each scalar input is expanded into a high-dimensional representation via a shared linear embedding layer:
    $X\leftarrow\texttt{Linear}(1\rightarrow D)$
    where $D$ is the embedding dimension.
  2. 2.
    Sparse Attention: The embedded inputs are passed through a NSA module. This module applies a combination of local attention (via sliding windows), block-wise compression, and block selection to reduce the quadratic complexity typically associated with full attention mechanisms.
  3. 3.
    TabMixer Module: Parallel to the attention mechanism, the embedded inputs are also processed through a TabMixer block, which applies channel- and token-wise MLP transformations with non-linear activations (e.g., GeLU and SiLU). This block captures complex interactions between features and across examples.
  4. 4.
    Fusion and Aggregation: The outputs from the attention and TabMixer modules are summed element-wise, and mean pooling is applied across the token (feature) dimension:
    $X\leftarrow\texttt{MeanPooling}(Y_{\text{attn}}+Z_{\text{mixer}})$
  5. 5.
    Prediction Head: A two-layer MLP with a GeLU activation reduces the feature dimension and produces the final classification logits corresponding to the number of output classes:
    $output=\texttt{MLP}(X)$
This architecture combines the global context modeling benefits of attention with the efficient token-feature mixing of MLP-based architectures, enabling accurate and scalable learning from tabular data while preserving both global and local context awareness.
4.3 TabNSA Architecture for Few Shot Learning
We propose TabNSA+Gemma, a hybrid model that integrates pretrained large language models with the TabNSA module to enhance tabular data classification. The architecture is designed to leverage the rich semantic representations of LLMs and adapt them to structured data classification tasks through attention-based modeling through the TabNSA module. This integration is aimed at enabling effective few-shot learning by utilizing the generalization capabilities of LLMs when labeled tabular data is limited.
4.3.1 TabNSA+Gemma
We leverage the instruction-tuned Gemma-1.1-2B model to generate meaningful representations from natural language prompts derived from tabular features. Specifically, our architecture integrates a pretrained lightweight Gemma language model (LLM) as a feature encoder with the specialized TabNSA module for downstream classification. The LLM encoder processes tokenized input sequences to produce contextualized token embeddings. Overall, the proposed architecture comprises three primary components: a partially frozen LLM encoder, a linear adapter layer, and the TabNSA model.
To balance computational efficiency and task-specific adaptation, we freeze most parameters of the Gemma encoder, selectively fine-tuning only the final two Transformer layers. This strategy preserves the general semantic knowledge acquired during pretraining while substantially reducing computational costs and memory usage. These embeddings undergo a dual-pooling mechanism, consisting of both max pooling and masked mean pooling, to derive compact representations. The max-pooled embedding captures salient features, while the mean-pooled embedding summarizes the overall context. A learnable linear adapter then projects the max-pooled embedding to match the dimensionality required by TabNSA, after which it is fused with the mean-pooled embedding via element-wise addition. This combined vector, encapsulating both prominent and summary information, serves as input to the TabNSA model for final classification. Hyperparameters used for TabNSA were determined through supervised learning experiments detailed in Section 5.4.
Let $\mathbf{H}\in\mathbb{R}^{B\times T\times D}$ denote the final hidden states from the LLM for a batch of $B$ samples, each with $T$ tokens length and embedding dimension $D$. Given an attention mask $\mathbf{M}\in\{0,1\}^{B\times T}$, we compute the pooled representations:
$\mathbf{h}_{\text{max}}$ $=\max_{t}\left(\mathbf{H}_{:,t,:}\odot\mathbf{M}_{:,t,:}+(1-\mathbf{M}_{:,t,:})\cdot(-\infty)\right)$ (12)
$\mathbf{h}_{\text{mean}}$ $=\frac{1}{\sum_{t=1}^{T}\mathbf{M}_{:,t}}\sum_{t=1}^{T}\mathbf{H}_{:,t,:}\odot\mathbf{M}_{:,t,:}$ (13)
where $\mathbf{M}_{:,t,:}=\mathbf{M}_{:,t}\mathbf{1}_{D}^{T}$ broadcasts the mask along the embedding dimension. These are combined via:
$\mathbf{h}_{\text{adapted}}=\mathbf{W}\mathbf{h}_{\text{max}}+\mathbf{b}+\mathbf{h}_{\text{mean}}$ (14)
with $\mathbf{W}\in\mathbb{R}^{D\times D}$, $\mathbf{b}\in\mathbb{R}^{D}$, yielding $\mathbf{h}_{\text{adapted}}\in\mathbb{R}^{B\times D}$ for TabNSA’s input dimension $D$. The final output is:
$\mathbf{y}=\text{TabNSA}(\mathbf{h}_{\text{adapted}})$ (15)
The resulting vector is then passed to the TabNSA model, which applies a combination of sparse attention and token-feature mixing to produce the final classification output. In our experiments, we adopt an $K$-shot training framework.
Figure 3: TabNSA+Gemma for Few-Shot Learning
4.3.2 Tabular-to-Text Prompt Transformation
To enable LLM-based few-shot learning on tabular inputs, we transform each tabular row into a structured natural language prompt. Specifically, for an input instance $\mathbf{x}=[x_{1},x_{2},...,$ $x_{n}]$ with $n$ features and label $y$, we format the prompt as:
"Feature_1: {$x_{1}$}, Feature_2: {$x_{2}$}, ..., Feature_n: {$x_{n}$}. What is the predicted label?"
For classification tasks, the model is expected to output one of the predefined label strings. During few-shot prompting, we prepend $k$ support examples (formatted similarly) to the query instance, e.g.,
[…]
Categorical features are mapped to human-readable strings, while numerical features are used verbatim, preserving original feature names to enhance interpretability. This prompt design helps the LLM leverage both structural and semantic information, aligning with recent best practices in LLM-based tabular learning.
Furthermore, we adapt the input token length dynamically between 512 and 2048 tokens depending on the specific dataset and the typical text length associated with each entity. This strategy balances computational efficiency and context coverage, effectively utilizing the full capacity of Gemma’s attention window.
5 Experiments and Results
5.1 Dataset
@@
In this work, we evaluated the performance of the TabNSA model and several baseline models across a diverse set of datasets. The assessment included 8 widely used binary classification datasets and 2 multi-class datasets sourced from the UCI repository, OpenML, and Kaggle. These datasets span a wide range of domains, including finance, business, chemistry, geography, ecology, image recognition, and sports, ensuring a comprehensive evaluation of the models. This diverse benchmark dataset collection enables us to systematically assess the performance across various domains and data structures. A detailed description of each dataset is provided in Table 2.
Experiments involving supervised learning and transfer learning were conducted on a local machine with an AMD Ryzen Threadripper PRO 5965WX 24-core CPU, 62 GB of RAM, and an NVIDIA RTX A4500 GPU with 20 GB of memory, as well as on P4V16 GPU nodes and NVIDIA H100-80 GPUs for the TabNSA+Gemma experiments.
@@
Table 2: Dataset details including abbreviation, number of classes, number of data points, and number of features.
@@
5.2 Preprocessing